# ACTIVE: Continuous Similarity Search for Vessel Trajectories

### v2

@@ 6.2. Evaluation on CSTS
Figure 6. Comparison of different methods on querying
6.2.1. Default
Figure 6(a) presents the query time for all methods on the two datasets under the default settings. CSTS demonstrates the best performance in terms of query time, attributed to its well-designed search algorithm with speed-up strategies. O-PRT has the longest search time, approximately 57 seconds on DK-AIS and 13 seconds on US-AIS, and its results are not displayed in the figure due to the significant disparity. H-TrajI and F-TrajI, which utilize an index based on entire trajectories, incur higher costs during the candidate search phase within the specified range. H-DFT and F-DFT perform relatively better than H/F-TrajI because the segment-based index offers increased efficiency during candidate searching. Both H/F-TrajI and H/F-DFT employ the Hausdorff or Fréchet distance measures, which increase the search time during distance computations considerably, as they consider all point pairs within the trajectories. TrajCL-TI and CLEAR-TI, the two AI-based methods, exhibit strong efficiency since they leverage GPU capabilities, enabling faster distance computations. O-PQT takes longer than CSTS, indicating that the SVTI in CSTS is more effective than the Quad-tree based index.
Figure 6(b) shows the hit rates of all methods on the two datasets under default settings. The hit rate of CSTS reaches 81% on DK-AIS and 87% on US-AIS, the highest among all methods. This superior performance is attributed to the proposed OTRD measure, which incorporates future movement trends of trajectories. O-PQT and O-PRT exhibit hit rates identical to that of CSTS, as they also employ the OTRD measure, which determines effectiveness in similar trajectory search. H-TrajI and H-DFT achieve the same hit rate because both use Hausdorff distance, while F-TrajI and F-DFT also perform equivalently in terms of hit rate due to their use of Fréchet distance. The Hausdorff and Fréchet measures show comparable hit rates but perform worse than OTRD because they rely on static or full-path comparisons, making them a poor fit for evolving trajectories. The two AI-based similarity computation methods, TrajCL-TI and CLEAR-TI, achieve the lowest hit rates because they are designed to cluster trajectory variants and are not optimized for computing similarity across all original trajectories.
@@ 6.2.1. Default
Figure 7. Effect of $l_{q}$
6.2.2. Effect of $l_{q}$
As shown in Figures 7(a) and 7(c), the query times for all methods increase as the query length $l_{q}$ increases. Among the traditional methods, CSTS exhibits the smallest rate of increase because it only evaluates the historical portion of trajectories, avoiding point-by-point checks for the entire trajectory. The query time for O-PQT also increases slightly since it uses the same similarity measure as CSTS, but its reliance on a point-based quad-tree yields higher search costs. H/F-TrajI and H/F-DFT exhibit a more pronounced growth in query time as $l_{q}$ increases, as each point in the historical part of the query trajectory is compared against the entire set of historical trajectories. In contrast, TrajCL-TI and CLEAR-TI show almost no variation in query time varying $l_{q}$, as these methods transform trajectories into fixed-length vectors, making them independent of query length. Figures 7(b) and 7(d) reveal no discernible trend in the hit rate, suggesting that the effect of $l_{q}$ on hit rate is relatively minor.
6.2.3. Effect of $r$
As shown in Figures 8(a) and 8(c), the query time for all methods increases with the range $r$, as a larger range results in more candidates to evaluate. H/F-TrajI and H/F-DFT experience the fastest growth in query time due to the higher computational cost of similarity computations for each candidate compared to OTRD and the AI-based methods. Among all methods, CSTS remains the most efficient, even as the range increases, owing to its optimized search algorithm. Figures 8(b) and 8(d) show that the hit rate decreases as the range increases. This is because a larger range includes more candidates. Some of these candidates might appear in the final similar trajectory results but not in the kkNN of $o$’s next position, leading to a reduced hit rate.
@@ 6.2.3. Effect of $r$
Figure 8. Effect of $r$
@@
Figure 9. Effect of $k$
6.2.4. Effect of $k$
As shown in Figure 9, increasing the value of $k$ leads to trends in query time and hit rate similar to those observed when increasing the value of $r$. Specifically, the query time increases, while the hit rate decreases. This occurs because the number of candidates is directly proportional to $k$, and a larger $k$ results in more candidates to evaluate, thereby increasing the time required for similarity computation. Moreover, the hit rate decreases as $k$ grows since $k$ serves as the denominator in the hit rate computation, making it more likely to produce a lower hit rate with larger $k$.
6.2.5. Effect of Data Size
We evaluate CSTS under varying data sizes, as shown in Figure 10. The settings for datasets D-1 through D-5 remain consistent with those described in Section B.1.2. As the data size increases, both the query time and the hit rate remain relatively stable. This is because the number of candidates remains consistent regardless of the dataset size. Consequently, the query time is primarily determined by the similarity computation rather than the candidate-finding phase, which incurs negligible overhead.
@@ 6.2.5. Effect of Data Size
Figure 10. Effect of data size
6.2.6. Effect of $\theta$ and $\alpha$
We examine the impact of the decay factor $\theta$ and the trade-off factor $\alpha$ on CSTS. Since these parameters do not influence efficiency, only their effects on the hit rate are reported. As shown in Figure 11(a), the hit rate decreases slightly as $\theta$ increases for both datasets. This is because a smaller $\theta$ results in a higher decay for earlier points in the historical portion of the query object. Consequently, points closer to the current position of the query object are given higher weight during similarity computation, leading to more accurate trajectory matches.
Similarly, Figure 11(b) shows a slight decline in the hit rate as $\alpha$ increases. This indicates that the Historical Trajectory Distance plays a more significant role in the similarity measure compared to the coarser-grained Target-Trajectory Distance.
@@ 6.2.6. Effect of $\theta$ and $\alpha$
Figure 11. Effect of $\theta$ and $\alpha$
7. Related Work
@@
Trajectory similarity measures. Trajectory similarity measures form the foundation of trajectory analysis. Dynamic Time Warping (Yi et al., 1998; Ying et al., 2016) allows for flexible alignment and comparison of time-series data, accommodating variations in speed and temporal shifts. Longest Common Subsequence (Vlachos et al., 2002) offers robustness to noise and spatiotemporal deformations by identifying the longest subsequence shared between trajectories. Additionally, Fréchet distance (Alt and Godau, 1995; De Berg et al., 2013; Agarwal et al., 2014) and edit distance-based measures (Chen and Ng, 2004; Chen et al., 2005; Ranu et al., 2015) provide nuanced assessments of trajectory similarity, considering geometric constraints and structural differences. The Hausdorff distance (Hangouet, 1995; Bai et al., 2011) measures how far the points of one set are from the points of the other set. These measures (Toohey and Duckham, 2015; Magdy et al., 2015) serve as the cornerstone for trajectory analysis. However, these measures fail to consider the significance of individual points, the trends within historical trajectories, and the future movement patterns of objects.
@@
Learning-based trajectory similarity computation. Recent advancements in trajectory similarity search have explored learning-based approaches (Li et al., 2018; Fu and Lee, 2020; Liu et al., 2020; Deng et al., 2022; Fang et al., 2022; Mao et al., 2022; Li et al., 2024a) to improve the accuracy and efficiency of similarity computations. These methods often leverage deep learning models to automatically learn representations of trajectories from raw spatiotemporal data. However, learning-based trajectory similarity computation has several limitations that make it less suitable for real-time, continuous vessel trajectory search. First, these methods often rely on down-sampling or distortion of trajectories, which leads to the identification of trajectories as similar, even when they lack true alignment with the original paths. This compromises their ability to accurately capture the underlying similarity between the original trajectories. Second, these methods typically focus on computing the similarity between two given trajectories rather than provide efficiently similarity search method on a large trajectory database. Third, they are not well-suited for real-time processing. The continuous updates in vessel trajectories would necessitate frequent model re-training or recalibration, imposing substantial computational overhead and significantly degrading performance. Finally, learning-based models are generally designed for offline batch processing, making them unsuitable for dynamic and real-time maritime navigation.
@@
Similar trajectory search. Studies on similar trajectory search (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023) and similarity join (Bakalov et al., 2005b, a; Ta et al., 2017; Shang et al., 2017) have proposed various techniques to identify similar movement patterns. Some approaches (Hwang et al., 2006, 2005; Xie et al., 2017; Yuan and Li, 2019; Han et al., 2021; He et al., 2022; Fang et al., 2022) are widely used in domains like vehicle tracking and urban traffic monitoring, particularly in road networks, where vehicle movements are constrained by structured paths, making it easier to organize and index trajectories based on fixed road segments, but not suitable for vessel trajectories. Furthermore, they mainly designed for existing trajectory similarity methods to compare entire trajectories, which is not suitable for real-time similar vessel trajectory search. These methods are less effective for real-time applications, where the trajectory evolves over time and must be analyzed incrementally. As a result, current trajectory similarity search methods are not well-suited to solve the problem of real-time, continuous vessel trajectory search.
8. Conclusion and Future Work
@@
Acknowledgements.
We acknowledge the support of the EU’s Horizon Europe research program under grant agreement No. 101070279 (MobiSpaces).
Appendix A Algorithm of Segment Partitioning
Algorithm 3 outlines the segment partitioning process in detail. Initially, a segment set $\mathcal{S}$ is initialized (line 1). For each trajectory $T_{i}$, the algorithm identifies the optimal segmentation (lines 2–16). Specifically, a list is created for $T_{i}$’s segment set, where the $j$th item represents the current optimal segment set up to point $p^{j}$ (line 3). Additionally, another list is initialized to store the total MBR area, where the $j$th item represents the minimum total MBR area up to point $p^{j}$ (line 4). For each point $p^{j}$ in $T_{i}$, the minimum total MBR area up to $p^{j}$ is computed (lines 5–15). If $j$ is less than $l_{\textit{min}}$, the area is set to 0; otherwise, it is initialized to $\infty$ (lines 6–7). For each possible starting point $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$, the MBR area of the segment $\langle p^{j-l+1},...,p^{j}\rangle$ is computed (lines 8–10). If the total MBR area is smaller than the current value, both the area and the segment set are updated (lines 11–15). Once all points in $T_{i}$ have been processed, the final segment set for $T_{i}$ is added to the overall segment set $\mathcal{S}$ (line 16). Ultimately, the algorithm returns the overall segment set $\mathcal{S}$ (line 17).
2:Output: A segment set $\mathcal{S}$.
3:Initialize a segment set $\mathcal{S}$ $\leftarrow$ $\emptyset$
4:for $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$ in $\mathcal{T}$ do
5:  Initialize a list for $T_{i}$’s segment set $S_{i}$ $\leftarrow$ $\emptyset$
6:  Initialize a list for the total MBR area $F$ $\leftarrow$ $\emptyset$
7:  for $j$ in $[1,n]$ do
8:   if $j\leq l_{\textit{min}}$ then $F[j]$ $\leftarrow$ 0
9:   else $F[j]$ $\leftarrow$ $\infty$
10:   for $l$ in $l_{\textit{min}},l_{\textit{max}}]$ do
11:     if $j-l\geq 0$ then
12:      $m_{l}$ $\leftarrow$ $\mathcal{M}_{area}$($p_{i}^{j-l+1}$, $p_{i}^{j}$)
13:      if jlj-l is 0 then $m$ $\leftarrow$ $m_{l}$
14:      else $m$ $\leftarrow$ $F[j-l]+m_{l}$
15:      if $F[j]>m$ then
16:        $F[j]$ $\leftarrow$ $m$
18:  $\mathcal{S}\leftarrow\mathcal{S}\cup S_{i}[n]$
19:return $\mathcal{S}$
Appendix B Evaluation
B.1. Evaluation on SVTI
B.1.1. Default
The performance of the different index methods under the default parameter settings is presented in Figure 12. In terms of construction time, SVTI takes 42 minutes to build on the DK-AIS dataset and 40 minutes on the US-AIS dataset. The index sizes of SVTI are 196 MB and 235 MB on DK-AIS and US-AIS, respectively. Among the five indices, PRT has the longest construction time, exceeding one day for both datasets due to its point-based design. Consequently, PRT also has the largest index size. Next, TrajI is the most efficient index to construct, as it is based on entire trajectories. This design minimizes insertion times, making its construction time directly proportional to the number of insertions. The original segmentation method for DFT, designed for road networks, cannot be applied directly to AIS data, so our segment partitioning method is used. Therefore, the construction time and index size of DFT are comparable to those of SVTI. Lastly, PQT exhibits relatively higher efficiency compared to PRT because the Quad-tree based PQT benefits from faster theoretical construction times compared to the R-tree based PRT.
@@ B.1.1. Default
Figure 12. Comparison of index construction time and size
B.1.2. Effect of Data Size.
We analyze the effect of varying data sizes on construction time and index size, with results shown in Figures 13(a) and 13(b), respectively. The datasets D-1, D-2, D-3, D-4, and D-5 contain approximately 43, 86, 130, 172, and 216 million points for DK-AIS, and 31, 62, 93, 123, and 154 million points for US-AIS. As the data size increases, the number of segments and corresponding insertions also grow. This leads to a proportional increase in both construction time and index size.
@@
Figure 13. Effect of data size
B.1.3. Effect of $l_{\textit{max}}$ and $l_{\text{min}}$

### v3

@@ V-B Evaluation on CSTS
V-B1 Default
Fig. 6(a) presents the query time for all methods on the two datasets under the default settings. CSTS demonstrates the best performance in terms of query time, attributed to its well-designed search algorithm with speed-up strategies. O-PRT has the longest search time, approximately 57 seconds on DK-AIS and 13 seconds on US-AIS, and its results are not displayed in the figure due to the significant disparity. H-TrajI and F-TrajI, which utilize an index based on entire trajectories, incur higher costs during the candidate search phase within the specified range. H-DFT and F-DFT perform relatively better than H/F-TrajI because the segment-based index offers increased efficiency during candidate searching. Both H/F-TrajI and H/F-DFT employ the Hausdorff or Fréchet distance measures, which increase the search time during distance computations considerably, as they consider all point pairs within the trajectories. TrajCL-TI and CLEAR-TI, the two AI-based methods, exhibit strong efficiency since they leverage GPU capabilities, enabling faster distance computations. O-PQT takes longer than CSTS, indicating that the SVTI in CSTS is more effective than the Quad-tree based index.
@@ V-B1 Default
Figure 6: Comparison of different methods on querying.
Fig. 6(b) shows the hit rates of all methods on the two datasets under default settings. The hit rate of CSTS reaches 81% on DK-AIS and 87% on US-AIS, the highest among all methods. This superior performance is attributed to the proposed OTRD measure, which incorporates future movement trends of trajectories. O-PQT and O-PRT exhibit hit rates identical to that of CSTS, as they also employ the OTRD measure, which determines effectiveness in similar trajectory search. H-TrajI and H-DFT achieve the same hit rate because both use Hausdorff distance, while F-TrajI and F-DFT also perform equivalently in terms of hit rate due to their use of Fréchet distance. The Hausdorff and Fréchet measures show comparable hit rates but perform worse than OTRD because they rely on static or full-path comparisons, making them a poor fit for evolving trajectories. The two AI-based similarity computation methods, TrajCL-TI and CLEAR-TI, achieve the lowest hit rates because they are designed to cluster trajectory variants and are not optimized for computing similarity across all original trajectories.
@@
Figure 7: Effect of $l_{q}$.
V-B2 Effect of $l_{q}$
As shown in Figs. 7(a) and 7(c), the query times for all methods increase as the query length $l_{q}$ increases. Among the traditional methods, CSTS exhibits the smallest rate of increase because it only evaluates the historical portion of trajectories, avoiding point-by-point checks for the entire trajectory. The query time for O-PQT also increases slightly since it uses the same similarity measure as CSTS, but its reliance on a point-based quad-tree yields higher search costs. H/F-TrajI and H/F-DFT exhibit a more pronounced growth in query time as $l_{q}$ increases, as each point in the historical part of the query trajectory is compared against the entire set of historical trajectories. In contrast, TrajCL-TI and CLEAR-TI show almost no variation in query time varying $l_{q}$, as these methods transform trajectories into fixed-length vectors, making them independent of query length. Figs. 7(b) and 7(d) reveal no discernible trend in the hit rate, suggesting that the effect of $l_{q}$ on hit rate is relatively minor.
@@ V-B2 Effect of $l_{q}$
Figure 8: Effect of $r$.
V-B3 Effect of $r$
As shown in Figs. 8(a) and 8(c), the query time for all methods increases with the range $r$, as a larger range results in more candidates to evaluate. H/F-TrajI and H/F-DFT experience the fastest growth in query time due to the higher computational cost of similarity computations for each candidate compared to OTRD and the AI-based methods. Among all methods, CSTS remains the most efficient, even as the range increases, owing to its optimized search algorithm. Figs. 8(b) and 8(d) show that the hit rate decreases as the range increases. This is because a larger range includes more candidates. Some of these candidates might appear in the final similar trajectory results but not in the kkNN of $o$’s next position, leading to a reduced hit rate.
@@ V-B3 Effect of $r$
Figure 9: Effect of $k$.
V-B4 Effect of $k$
As shown in Fig. 9, increasing the value of $k$ leads to trends in query time and hit rate similar to those observed when increasing the value of $r$. Specifically, the query time increases, while the hit rate decreases. This occurs because the number of candidates is directly proportional to $k$, and a larger $k$ results in more candidates to evaluate, thereby increasing the time required for similarity computation. Moreover, the hit rate decreases as $k$ grows since $k$ serves as the denominator in the hit rate computation, making it more likely to produce a lower hit rate with larger $k$.
V-B5 Effect of Data Size
We evaluate CSTS under varying data sizes, as shown in Fig. 10. The settings for datasets D-1 through D-5 remain consistent with those described in Section A-A3. As the data size increases, both the query time and the hit rate remain relatively stable. This is because the number of candidates remains consistent regardless of the dataset size. Consequently, the query time is primarily determined by the similarity computation rather than the candidate-finding phase, which incurs negligible overhead.
@@ V-B5 Effect of Data Size
Figure 10: Effect of data size.
V-B6 Effect of $\theta$ and $\alpha$
We examine the impact of the decay factor $\theta$ and the trade-off factor $\alpha$ on CSTS. Since these parameters do not influence efficiency, only their effects on the hit rate are reported. As shown in Fig. 11(a), the hit rate decreases slightly as $\theta$ increases for both datasets. This is because a smaller $\theta$ results in a higher decay for earlier points in the historical portion of the query object. Consequently, points closer to the current position of the query object are given higher weight during similarity computation, leading to more accurate trajectory matches.
Similarly, Fig. 11(b) shows a slight decline in the hit rate as $\alpha$ increases. This indicates that the Historical Trajectory Distance plays a more significant role in the similarity measure compared to the coarser-grained Target-Trajectory Distance.
@@ V-B6 Effect of $\theta$ and $\alpha$
Figure 11: Effect of $\theta$ and $\alpha$.
Figure 12: Effect of speed-up strategies.
V-C Ablation Study
By default, CSTS incorporates segment pruning (Strategy 1), $k$-bound pruning (Strategy 2), and the incremental rule (Strategy 3). These three strategies are solely designed to improve efficiency and do not reduce the hit rate. In Fig. 12, we report their impact on query time. Among the strategies, the incremental rule (S3) proves to be the most effective, reducing the query time by at least threefold. This is the evidence of the efficiency of the incremental rule in optimizing CSTS. The second-most effective strategy is $k$-bound pruning (S2), which shows comparable performance to S3. Segment pruning (S1), however, provides relatively less improvement on both datasets. This is because it introduces additional computational costs to compute the distance between a point and the minimum bounding rectangle of a segment.
(a) Time vs. $g$
(b) Hit vs. $g$
Figure 13: Effect of granularity.
The granularity adjustment (Strategy 4) is not applied in CSTS by default, as it may introduce errors. To evaluate the effectiveness of this strategy, we vary the granularity parameter $g$. Referring to Fig. 13, both query time and hit rate decrease as $g$ increases. With a larger $g$, more points are disregarded during similarity computation, resulting in reduced query time. However, the hit rate declines because ignoring points introduces errors.
Figure 14: Case study.
V-D Case Study on Similarity Measures
Fig. 14 presents case studies on DK-AIS and US-AIS, illustrating the differences in results when applying different similarity measures. The orange line represents a moving object $o$, including both its current trajectory and the projected destination. The black line indicates the most similar trajectory identified using the OTRD measure. In comparison with other similarity measures such as Hausdorff, Fréchet, TrajCL, and CLEAR, the OTRD measure demonstrates a clear advantage. The other four measures focus predominantly on aligning with $o$’s current trajectory, resulting in matches that are less aligned with $o$’s overall path. Conversely, the trajectory identified by OTRD not only aligns closely with $o$’s current trajectory but also incorporates the future trend and destination, offering a more holistic similarity result. This highlights the strength of OTRD at capturing the dynamic and predictive aspects of trajectory comparison, making it suitable for applications where future trends and endpoints are important.
VI Related Work
@@
Trajectory similarity measures. Trajectory similarity measures form the foundation of trajectory analysis. Dynamic Time Warping [17, 18] allows for flexible alignment and comparison of time-series data, accommodating variations. Longest Common Subsequence [19] offers robustness to noise and spatiotemporal deformations by identifying the longest subsequence shared between trajectories. Additionally, Fréchet distance [20, 21, 22] and edit distance-based measures [25, 26, 27] provide nuanced assessments of trajectory similarity, considering geometric constraints and structural differences. The Hausdorff distance [23, 24] measures how far the points of one set are from the points of the other set. These measures [30, 31] serve as the cornerstone for trajectory analysis. However, these measures fail to consider the significance of individual points, the trends within historical trajectories, and the future movement patterns of objects.
@@
Learning-based trajectory similarity computation. Recent advancements in trajectory similarity search have explored learning-based approaches [32, 33, 34, 35, 36, 37, 29] to improve the accuracy and efficiency of similarity computations. These methods often leverage deep learning models to learn representations of trajectories from raw spatiotemporal data. However, this method has several limitations that make it less suitable for real-time, continuous vessel trajectory search. First, these methods often rely on down-sampling or distortion of trajectories, which leads to the identification of trajectories as similar, even when they lack true alignment with the original paths. This compromises their ability to accurately capture the underlying similarity between the original trajectories. Second, these methods typically focus on computing the similarity between two given trajectories rather than provide efficiently similarity search method on a large trajectory database. Third, they are not well-suited for real-time processing. The continuous updates in vessel trajectories would necessitate frequent model re-training or recalibration, imposing substantial computational overhead and significantly degrading performance. Finally, learning-based models are generally designed for offline batch processing, making them unsuitable for dynamic and real-time maritime navigation.
@@
Similar trajectory search. Studies on similar trajectory search [7, 8, 9, 10, 11, 12] and similarity join [13, 14, 15, 16] have proposed various techniques to identify similar movement patterns. Some approaches [38, 39, 8, 40, 41, 11, 36] are widely used in domains like vehicle tracking and urban traffic monitoring, particularly in road networks, where vehicle movements are constrained by structured paths, making it easier to organize and index trajectories based on road segments, but not suitable for vessel trajectories. Furthermore, they mainly designed for existing trajectory similarity methods to compare entire trajectories, which is not suitable for real-time similar vessel trajectory search. These methods are less effective for real-time applications, where the trajectory evolves over time and must be analyzed incrementally. As a result, current trajectory similarity search methods are not well-suited to solve the problem of real-time, continuous vessel trajectory search.
VII Conclusion and Future Work
@@
Acknowledgments
We acknowledge the support of European Union’s funded Project MobiSpaces under grant agreement no 101070279.
Appendix A Evaluation
A-A Evaluation on SVTI
A-A1 Settings
We evaluate SVTI and four traditional indexing techniques—PQT, PRT, TrajI, and DFT—in terms of index construction time and index size. To assess the performance of SVTI, we vary three parameters with the following settings (default values in bold):
  • $l_{\textit{min}}$: The minimum segment lengths. Setting: 20, 25, 30, 35, 40.
  • $l_{\textit{max}}$: The maximum segment lengths. Setting: 40, 45, 50, 55, 60.
A-A2 Default
The performance of the different index methods under the default parameter settings is shown in Fig. 15. In terms of construction time, SVTI takes 42 minutes to build on the DK-AIS dataset and 40 minutes on the US-AIS dataset. The index sizes of SVTI are 196 MB and 235 MB on DK-AIS and US-AIS, respectively. Among the five indices, PRT has the longest construction time, exceeding one day for both datasets due to its point-based design. Consequently, PRT also has the largest index size. Next, TrajI is the most efficient index to construct, as it is based on entire trajectories. This design minimizes insertion times, making its construction time directly proportional to the number of insertions. The original segmentation method for DFT, designed for road networks, cannot be applied directly to AIS data, so our segment partitioning method is used. Therefore, the construction time and index size of DFT are comparable to those of SVTI. Lastly, PQT exhibits relatively higher efficiency compared to PRT because the Quad-tree based PQT benefits from faster theoretical construction times compared to the R-tree based PRT.
@@ A-A2 Default
Figure 15: Comparison of index construction time and size.
A-A3 Effect of Data Size.
We analyze the effect of varying data sizes on construction time and index size, with results shown in Figs. 16(a) and 16(b), respectively. The datasets D-1, D-2, D-3, D-4, and D-5 contain approximately 43, 86, 130, 172, and 216 million points for DK-AIS, and 31, 62, 93, 123, and 154 million points for US-AIS. As the data size increases, the number of segments and corresponding insertions also grow. This leads to a proportional increase in both construction time and index size.
@@
Figure 16: Effect of data size.
A-A4 Effect of $l_{\textit{max}}$ and $l_{\text{min}}$